# thanks: corresponding author

### v3

@@ Appendix E Proof of Theorems 3 and 4
$=$ $\max_{\mbox{\tiny$\begin{array}[]{c}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}},\bm{\rho}=\bm{\rho}^{\dagger},\\ \text{rank}(\bm{\rho})=(2r_{1},\dots,2r_{n-1}),\\ \operatorname{trace}(\bm{\rho})=0,\|\bm{\rho}\|_{F}\leq\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}\end{array}$}}\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$
@@
$\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle$ (56)
$=$ $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle^{2}-\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle$
@@
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ (59)
$\leq$ $(1+b)\big{(}\big{(}1-\frac{d^{n+1}(1-\delta)\mu}{K(d^{n}+1)}+\frac{d^{2n+1}(1+\delta)^{2}\mu^{2}}{K^{2}(d^{n}+1)^{2}}\big{)}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big{)}$
@@
where we define the error term $e=\frac{2\mu c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{2\mu^{2}c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M}$, and $c_{3},c_{4}$ are positive constants. The last two lines follow assumptions with regard to the step size and initialization: $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and $b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}$ which guarantee that the convergence rate satisfies
@@
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}=\|\mathcal{P}_{\operatorname{trace}}(\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$ (62)
@@ Appendix F Proof of Theorem 5
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}$ $=$ $\bigg{\|}\mathcal{P}_{\operatorname{trace}}\bigg{(}\text{SVD}_{\bm{r}}^{tt}\bigg{(}\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg{)}-\bm{\rho}^{\star}\bigg{)}\bigg{\|}_{F,2\overline{r}}$ (63)
@@
$=$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}(1+\delta)(\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))-\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle\leq\delta\|\bm{\rho}^{\star}\|_{F}.$
@@ Appendix G Auxiliary Materials
${\bm{A}}_{1}{\bm{A}}_{2}\cdots{\bm{A}}_{N}-{\bm{A}}_{1}^{\star}{\bm{A}}_{2}^{\star}\cdots{\bm{A}}_{N}^{\star}=\sum_{i=1}^{N}\bm{A}_{1}^{\star}\cdots\bm{A}_{i-1}^{\star}(\bm{A}_{i}-\bm{A}_{i}^{\star})\bm{A}_{i+1}\cdots\bm{A}_{N}.$ (66)

### v4

@@ Appendix E Proof of Theorems 3 and 4
$=$ $\hskip-42.67912pt\max_{\mbox{\tiny$\begin{array}[]{c}\bm{\rho}\in\mathbb{C}^{d^{n}\times d^{n}},\bm{\rho}=\bm{\rho}^{\dagger},\\ \text{rank}(\bm{\rho})=(2r_{1},\dots,2r_{n-1}),\\ \operatorname{trace}(\bm{\rho})=0,\|\bm{\rho}\|_{F}\leq\|\bm{\rho}_{1}-\bm{\rho}_{2}\|_{F}\end{array}$}}\hskip-28.45274pt\langle\bm{\rho}_{1}-\bm{\rho}_{2},\bm{\rho}\rangle.$
@@
$\langle\bm{\rho}^{(\tau)}-\bm{\rho}^{\star},\nabla g(\bm{\rho}^{(\tau)})\rangle$ $\!\!\!\!=\!\!\!\!$ $\sum_{k=1}^{K}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle^{2}-\sum_{k=1}^{K}\eta_{k}\langle\bm{A}_{k},\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\rangle$ (56)
@@
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$ $\!\!\!\!\leq\!\!\!\!$ $(1+b)\big(\big(1-\frac{d^{n+1}(1-\delta)\mu}{K(d^{n}+1)}+\frac{d^{2n+1}(1+\delta)^{2}\mu^{2}}{K^{2}(d^{n}+1)^{2}}\big)\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}^{2}+e\big)$ (59)
@@
where we define the error term
$e=\frac{2\mu c_{1}d\overline{r}\sqrt{nd^{n}(1+\delta)\gamma_{t}(\bm{\rho}^{\star})\log n}}{K\sqrt{(d^{n}+1)M}}\|\bm{\rho}^{(\tau)}-\bm{\rho}^{\star}\|_{F}+\frac{2\mu^{2}c_{1}^{2}(1+\delta)\overline{r}^{2}nd^{n+2}\gamma_{t}(\bm{\rho}^{\star})\log n}{K^{2}(d^{n}+1)M},$ (60)
and $c_{3},c_{4}$ are positive constants. The last two lines follow assumptions with regard to the step size and initialization: $\frac{bK(d^{n}+1)}{d^{n}(1+b)(1-\delta)}<\mu\leq\frac{(d-1)(d^{n}+1)(1-\delta)K}{(1+\delta)^{2}d^{n+1}}$ and $b<\frac{(d-1)(1-\delta)^{2}}{1+\delta^{2}+(4d-2)\delta}$ which guarantee that the convergence rate satisfies
@@
$\|\bm{\rho}^{(\tau+1)}-\bm{\rho}^{\star}\|_{F}^{2}$
$\!\!\!\!=\!\!\!\!$ $\|\mathcal{P}_{\operatorname{trace}}(\text{SVD}_{\bm{r}}^{tt}(\bm{\rho}^{(\tau)}-\mu\nabla g(\bm{\rho}^{(\tau)})))-\bm{\rho}^{\star}\|_{F,2\overline{r}}^{2}$
@@ Appendix F Proof of Theorem 5
$\|\bm{\rho}^{(0)}-\bm{\rho}^{\star}\|_{F}$
$\!\!\!\!=\!\!\!\!$ $\bigg\|\mathcal{P}_{\operatorname{trace}}\bigg(\text{SVD}_{\bm{r}}^{tt}\bigg(\sum_{k=1}^{K}\frac{K(d^{n}+1)}{d^{n}}(\langle\bm{A}_{k},\bm{\rho}^{\star}\rangle+\eta_{k})\bm{A}_{k}\bigg)-\bm{\rho}^{\star}\bigg)\bigg\|_{F,2\overline{r}}$
@@
$\!\!\!\!=\!\!\!\!$ $\max_{\widetilde{\bm{\rho}}\in\overline{\mathbb{X}}_{2\bm{r}}}(1+\delta)(\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle+\operatorname{trace}(\bm{\rho}^{\star})\operatorname{trace}(\widetilde{\bm{\rho}}))-\langle\bm{\rho}^{\star},\widetilde{\bm{\rho}}\rangle$
$\!\!\!\!\leq\!\!\!\!$ $\delta\|\bm{\rho}^{\star}\|_{F}.$
@@ Appendix G Auxiliary Materials
${\bm{A}}_{1}{\bm{A}}_{2}\cdots{\bm{A}}_{N}-{\bm{A}}_{1}^{\star}{\bm{A}}_{2}^{\star}\cdots{\bm{A}}_{N}^{\star}$ (67)
$\!\!\!\!=\!\!\!\!$ $\sum_{i=1}^{N}\bm{A}_{1}^{\star}\cdots\bm{A}_{i-1}^{\star}(\bm{A}_{i}-\bm{A}_{i}^{\star})\bm{A}_{i+1}\cdots\bm{A}_{N}.$